# Bayesian Optimization in Language Space: An Eval-Efficient AI Self-Improvement Framework

### v2

@@ Abstract
Large Language Models (LLMs) have recently enabled self-improving AI, i.e., AI that iteratively generates, evaluates, and refines its own outcomes. Recent studies have shown that self-improving AI focusing on prompt optimization (i.e., context engineering) can outperform state-of-the-art reinforcement learning fine-tuned LLMs. Here, their ‘performance’ is typically measured by query efficiency—the number of LLM-generated solution samples required to reach a specified performance threshold. However, in many societal applications, the primary limitation is not generating new solutions but evaluating them. For instance, evaluating an ad’s effectiveness requires significant human feedback, which is far more costly and time-consuming than generating a candidate ad. To optimize for the evaluation efficiency objective, a natural approach is to extend Bayesian Optimization (BO), a framework proven optimal for evaluation efficiency, to the language domain. However, the difficulty of directly estimating suitable acquisition functions in LLMs’ minds makes this extension challenging. This paper overcomes this challenge by proving that the combination of the simple and widely used Best-of-N selection strategy and simple textual gradients (i.e., textual edits from a critic model), statistically emulates the behavior of the gradients on the canonical Upper Confidence Bound (UCB) acquisition function, which induces optimal exploration in terms of evaluation efficiency. Based on this result, we propose TextGrad-Best-of-N Bayesian Optimization (T-BoN BO), a simple and eval-efficient language-space Bayesian optimization framework for AI self-improvement. We also empirically validate T-BoN BO by applying it to automated ad alignment tasks for persona distribution, demonstrating its superior performance compared to popular state-of-the-art baselines. In sum, T-BoN BO is a simple but principled framework for AI self-improving system design that bridges prompt optimization with classical Bayesian optimization.
Keywords: Self-improving AI, Evaluation efficiency, Bayesian optimization, Context Engineering
@@ 1 Introduction
1.1 Self-improving AI
Data-centric decision making has long relied on a human-driven cycle of historical data analysis, solution proposal, and experimentation. This human-driven iterative approach has supported a wide range of business decisions, ranging from website design (hauser2009website) to product recommendations (nandy2021b). We briefly describe this paradigm using the example of ad-optimization; see the left panel of Figure 1. The process starts with generating a set of ad creatives based on a creative brief. This task is often done by ad agencies or marketers, who rely on their understanding of the product, theories of consumer behavior, market conditions, and feedback from prior ad campaigns. This ad generation process typically takes weeks, and serves as the main bottleneck in the iterative ad-optimization process. Once we have a set of candidate ads, the manager moves to evaluation, i.e., feeds these ads into advertising platforms or consumer surveys to identify the best performing one, and this step usually takes a few days. Finally, to complete this feedback loop, the manager analyzes the performance of the ads from this campaign and uses the insights learned to generate a new slate of candidate ads; and the iterative process continues.
Although effective in many settings, this human-driven workflow suffers from drawbacks. First, the speed with which firms can iterate over this process depends critically on how quickly humans can generate new candidate solutions. Historically, generating new candidate solutions (or ads in the case of our example) has been time-consuming, and this step formed a key bottleneck in speeding up the process. Further, there exists no formal framework for how human agents can analyze and learn from the performance of prior solutions and leverage them to propose better candidates, and this step is often left to managerial intuition. Indeed, much of the research and managerial focus on data-driven business solutions has largely centered on the evaluation step, i.e., given a set of candidate solutions/options, how can we identify the most effective one – and solutions such as A/B tests and Multi-Armed Bandits (MAB) experiments have been proposed and tested (kohavi2020trustworthy; fiez_etal_2024).
Ad generation (Step 1)(Weeks)Analysis (Step 3)(Days)Ad evaluation (Step 2)(Days-weeks)HumanBottleneckAd generation (Step 1)(seconds)Analysis (Step 3)(minutes–hours)Ad evaluation (Step 2)(Days-weeks)AIBottleneck
@@ 1.1 Self-improving AI
Recent advances in generative AI have begun to transform this human-driven paradigm by providing potential solutions to the issues discussed above. First, candidate generation is fast and scalable with generative AI models. For example, in the case of digital ads, generative models can produce high-quality creative outputs, including ad copy, images, and complete campaign concepts that adhere to brand guidelines in a matter of minutes (jansen2024automated; hartmann2025power). [^1] Indeed, a substantial share of creatives is now generated by AI, reflecting improvements in both quality and speed (coffee2025aiadbuying; Majic2025VibeMarketing). Second, Large language Models (LLMs) and related generative systems can now conduct sophisticated data analysis, summarizing performance patterns and identifying potential drivers of success (ferber2024context; guo2024ds; ghosh2024exploring; wiedemer2025videomodelszeroshotlearners), providing a pathway for analyzing historical data and improving candidate solutions. As such, with the new capabilities enabled by AI, the time required for previous data analysis and generation of new candidate solutions are significantly reduced.
Given these developments, there is immense interest in the development and deployment of AI-based automation of the complete generation–evaluation-analysis loop at scale, allowing systems to update themselves continuously in response to emerging data (See right panel of Figure 1). In this loop, generative models produce new solution candidates, data-science agents conduct evaluation of the solution candidates, and multi-modal reasoning LLM-based analytical agents interpret the evaluation results and synthesize insights to guide the next generation step. This type of automated cycle running system is often referred to as the self-improving AI or self-evolving AI (MantiaChatterjeeLee2025; chen2025xbench; silver2025welcome).
This transformation has an important implication for practice: it dramatically lowers the marginal cost of iteration, making continuous experimentation feasible not only for large firms or organizations with dedicated analytics teams but also for small and medium-sized firms that historically lacked the creative and analytical capacity to run many cycles in parallel. When generation and analysis become cheap and largely automated, even resource-constrained organizations can maintain always-on optimization loops that adapt campaigns to evolving market conditions and consumer responses.
As generation and analysis become increasingly automated and inexpensive in this self-improving loop, the bottleneck in data-centric decision making shifts from proposing and interpreting candidates to evaluating them in the real world. Even if analysis can be performed quickly and new candidate solutions can be generated almost instantaneously, each iteration still has to be validated by interacting with the real environment and collecting outcome data. The effective pace of the cycle is therefore governed by evaluation-related constraints such as available traffic, experimentation budgets, and the amount of data required to obtain statistically reliable comparisons among candidates, rather than by generation or analysis speed.
For instance, in digital advertising, even when generating ad creatives via LLM queries is super fast, but evaluating an ad’s effectiveness still requires significant human feedback or large-scale experimentation (johnson2023inferno). A similar pattern arises in other scientific and engineering domains: protein structures or sequences can be proposed rapidly by models such as AlphaFold (jumper2021highly), but their functional properties must still be confirmed through slow and noisy wet-lab assays; and chip layouts can be generated in silico by systems such as AlphaChip (goldie2024chip), yet rigorous evaluation depends on physical fabrication and benchmarking. As generation and analysis become increasingly automated and inexpensive, the evaluation step thus emerges as the primary bottleneck that limits the overall cadence of self-improvement. As such, any societal and business AI system we design needs to be evaluation-efficient, i.e., reach a high-performing solution using as few costly evaluations as possible, rather than optimizing for a few inexpensive generations.
1.2 Prompt optimization-based self-improving AI
A natural question is, then, how the computer science literature has instantiated self-improving AI in concrete algorithmic frameworks. To date, a central line of success in self-improving AI has arisen in the domain of iterative prompt optimization, where AI agents iteratively refine their own prompts in response to performance feedback. agrawal2025gepa showed that iterative prompt optimization-based self-improving AI can be much more efficient than state-of-the-art model fine-tuning techniques such as GRPO (liu2024deepseek). Specifically, given any limited budget for the number of LLM generation query calls, their proposed iterative prompt optimization method, named GEPA (Genetic-Pareto), outperformed GRPO across their experiments. Those experiments included benchmarks such as complex questions across multiple Wikipedia articles, strict instruction-following under constraints, retrieval-augmented fact verification to support or refute claims, and maintaining answer quality while avoiding privacy leakage.
In all those benchmarks, the primary efficiency metric was generation efficiency. i.e., the number of LLM query calls required to generate candidate answers to reach a target performance level. Likewise, prior literature on prompt-based self-improving AI (agrawal2025gepa; zhang2025agentic; wang2025maestro) has focused on measuring efficiency in terms of generation efficiency, not evaluation efficiency. Indeed, leading approaches are often specified as procedural heuristic search routines which finds near-optimal solution with small computations, e.g., population-based or random-search-style algorithms that treat prompts as individuals in an evolutionary process or as configurations in a heuristic pipeline. Specifically, they often maintain a pool or batch of prompts, repeatedly generate new variants via LLM-driven mutation, crossover, reflection, or trace-based resampling, score these candidates on held-out data or rollouts, retain the top-performing or Pareto-optimal prompts, and iterate this generate–score–select–mutate loop under a budget of LLM forward passes (10.5555/3692070.3692611; guo2024connecting; agrawal2025gepa; better_together). This mirrors the design philosophy of classical evolutionary and heuristic optimization, where the quality of the best solution is assessed as a function of the computation budget rather than the evaluation budget (nguyen2016understanding; ding2023quality; hu2024automated; liu2024evolution). Consequently, evaluation efficiency objective has been largely overlooked in the computer science literature.
However, as we discussed earlier, in self-improving AI for real-world societal and business problems, the primary bottleneck is no longer generating new solutions efficiently (generation efficiency) but evaluating them efficiently (evaluation efficiency). This shift in the dominant constraint calls for rethinking prompt-optimization-based self-improving AI frameworks so that they explicitly treat evaluation budget as the scarce resource to be allocated and optimized over.
@@ 1.3 Research agenda and Challenges
Motivated by the effectiveness of iterative prompt-optimization-based self-improving AI and the overlooked importance of evaluation efficiency in existing frameworks, we now turn to our central question: how can we design a prompt-optimization-based self-improving AI framework that is evaluation-efficient? In addition to evaluation efficiency, we would like the framework to be easy to implement, theoretically grounded, and empirically perform well. However, realizing these desiderata in a unified framework is nontrivial.
The key insight in this paper is that prompt-optimization-based self-improving AI follows a familiar pattern we saw in the optimization framework called Bayesian Optimization (BO) (frazier2018tutorial; srinivas2012information; papenmeier2025understanding). This insight is grounded in three observations: First, LLMs learn in-context from previous evaluations (brown2020language; coda2023meta; lu2024emergent; cahyawijaya2024llms), acting as Bayesian learners (xie2021explanation; wang2023large; wakayama2025context). Second, we iteratively choose a candidate prompt that generates a candidate solution, observe a noisy performance score associated with this solution, and then decide where to query next. Third, we have to be cognizant of three concerns: (1) we have no closed-form model linking prompts to scores, (2) we need to balance exploration and exploitation when choosing the next candidate solution, and (3) the goal is to optimize for the number of evaluations rather than candidate generation.
BO is a principled framework for tackling such black-box optimization problems with such features. In BO, we build a function called the acquisition function and choose the point with the top of the acquisition function as the next point to evaluate. Among many variants of BO, Upper Confidence Bound (UCB)-BO (srinivas2012information), a version of BO with UCB acquisition function, is proven to be evaluation-efficient, in the sense that no algorithm can achieve better regret in terms of both simple regret and expected cumulative regret (whitehouse2023sublinear; wang2023regret). Figure 2 illustrates the procedure of UCB-BO.
Figure 2: An example of a non-parallel UCB-BO procedure for four optimization steps. At each step, we choose the maximum (or multiple local maxima in the parallel UCB-BO) of the blue UCB function as the next point to evaluate.
As you can see from Figure 2, the UCB acquisition function is often non-concave. To resolve this issue, in practice, we often use parallel gradient-based UCB-BO (wilson2018maximizing), where we 1) start from multiple initial points and take gradients to find a set of local optima, and 2) choose those local optima as the points to evaluate next. Parallel gradient-based BO has shown surprising empirical effectiveness for high-dimensional problems (papenmeier2025understanding).
Building on the key insight, we propose to achieve evaluation efficiency by casting iterative prompt optimization-based self-improving AI as parallel gradient-based UCB Bayesian optimization in language space. At each iteration, we treat the prompts from the previous round as initialization points and take gradient steps to ascend the UCB acquisition function toward its local optima. We then select which prompts to evaluate from among these local optima. After each evaluation round, the LLM and its associated critic models implicitly update a surrogate belief over prompt quality via in-context learning and corresponding UCB, analogous to the posterior update in classical Bayesian optimization.
However, there are two challenges we need to overcome.
  • $\bullet$
    Challenge 1– Lack of gradient functions
    In prompt spaces, the notion of a gradient is ill-defined because they lack the inner-product structure of a Hilbert space required to measure the directions and magnitudes of change (young1988introduction). Thus, we first need to define a meaningful notion of a “gradient” in the domain of text. Although one might consider going around this challenge by embedding the prompt first, taking the gradient there, and reverting the prompt, this strategy runs into the well-known soft-to-discrete projection problem, where such reverted prompts break any clean correspondence between infinitesimal moves in embedding space and interpretable edits of the underlying natural-language prompt (cui2025automatic).
  • $\bullet$
    Challenge 2 – Taking gradients for UCB
@@
    Second, even if we can define meaningful gradients in prompt space, we do not have direct access to the underlying UCB acquisition function whose gradient we would like to follow. The surrogate posterior over prompt quality that induces UCB lives implicitly in the LLM, rather than as an explicit Gaussian-process-style model that exposes closed-form posterior means, variances, and their derivatives. Consequently, standard gradient-based BO techniques that differentiate the UCB constructed from a known surrogate cannot be applied naively: we must design a mechanism that elicits UCB-like, optimism-in-the-face-of-uncertainty behavior using only black-box text interactions and critic feedback, without ever computing the UCB function or its gradient in closed form.
In this paper, our goal is to address the above challenges and build an evaluation-efficient self-improving AI that optimally balances exploration and exploitation in the language space.
1.4 Proposed Method
Figure 3: Gradient steps of parallel gradient-based BO (Left image) and those of T-BoN BO (Right image). T-BoN BO extends parallel gradient-based BO to the language space.
@@ 1.4 Proposed Method
We propose a simple, evaluation-efficient iterative prompt-optimization framework, called TextGrad-Best-of-N Bayesian Optimization (T-BoN BO), that adapts parallel gradient-based UCB-BO for prompt optimization based self-improving AI (Figure 3), addressing both of the challenges discussed earlier.
To solve the first challenge, T-BoN BO builds on the recently proposed TextGrad (textgrad) framework that adapts gradient descent to discrete text by asking an external auxiliary LLM critic to propose targeted edits that improve a candidate; applying these edits induces a local, directional update, i.e., a “textual gradient”, in language space. This makes gradient-style search feasible without continuous relaxations or mapping the LLM prompts to the embedding space. For instance, consider an ad creative generation prompt for an image model: “A simple photo of a plant-based burger on a white plate.” A critic LLM might suggest: “Make the burger appear delicious, and show it being enjoyed at a barbecue.” Incorporating this edit yields: “A photorealistic image of a juicy plant-based burger with grill marks, served at a vibrant summer barbecue with friends.” This change is analogous to following a gradient direction—small, constructive, and aligned with improving performance according to the evaluation signal. textgrad show that textual gradients based on their approach enable consistent improvements across diverse scientific domains such as code-completion benchmarks, improving accuracy on graduate-level QA datasets, optimizing drug-like molecules, and refining radiotherapy planning for clinical objectives. These results establish textual gradients as an effective surrogate for numerical gradients when optimizing systems that operate in language space.
To address the second challenge, T-BoN BO chooses the prompt to evaluate next via a series of textual gradients which are selected under the Best-of-N principle (snell2024scaling), resulting in the gradient direction we call Best-of-N gradient. Moving in the textual gradient direction suggested by the critic exploits the expected improvement implied by past results but ignores uncertainty in those expectations. Therefore, it does not climb the UCB acquisition function. Best-of-N gradient bakes in exploration into the choice of gradient direction by adopting an optimism-based rule in the spirit of UCB’s optimism-in-the-face-of-uncertainty principle: it samples multiple local textual edits and moves along the direction the LLM judges most promising, i.e., the Best-of-$N$ candidate. That is, the Best-of-$N$ gradient is designed to simulate the implicit UCB acquisition gradient in language space.
@@
We now briefly describe our T-BoN BO algorithm. At each optimization iteration, in Step 1 (meta-reflection phase), we use the critic model to analyze the history of previously evaluated solutions and their scores, and update a meta-reflection that summarizes what has tended to work well or poorly. In Step 2 (Best-of-$N$ textual-gradient phase), we perform a sequence of Best-of-N gradient updates from the current prompt. That is, for each gradient step, (i) we propose $N$ textual-gradient edits that transform the current prompt into $N$ candidate prompts, (ii) each candidate prompt is passed to a solution-generation model to produce a corresponding candidate solution, (iii) conditioned on the meta-reflections, the critic model predicts which candidate solution is most likely to perform best, and (iv) we set the prompt associated with this predicted-best solution as the new current prompt. After iterating Step 2, we obtain a refined prompt and its associated solution to evaluate. In Step 3 (evaluation phase), we evaluate this solution to get a scalar score, which T-BoN BO then uses to update its history and meta-reflection for subsequent optimization steps.
1.5 Theoretical Guarantees and Empirical Evaluation
@@
We provide theoretical guarantees on the performance of our approach. We formally show that, under mild regularity conditions, the Best-of-$N$ gradient over locally sampled textual edits induces, in probability, ascent directions of the UCB acquisition function with an exploration parameter $\beta_{N}=\Theta(\sqrt{\ln N})$ (Theorem 2). In other words, T-BoN BO emulates parallel gradient-based UCB Bayesian Optimization in the implicit embedding space induced by the LLM, even though it operates purely in language space and does not construct an explicit surrogate or uncertainty model. Since UCB-BO is known to achieve sublinear regret and optimal evaluation-efficiency guarantees, T-BoN BO inherits these theoretical properties up to constant factors and approximation error stemming from textual gradients. In sum, we formally establish that T-BoN BO is theoretically equivalent to the parallel gradient-based implementation of Bayesian Optimization with the Upper Confidence Bound (UCB) acquisition function (srinivas2012information; wilson2018maximizing), which is proven to be evaluation-efficient (whitehouse2023sublinear; wang2023regret). This equivalence establishes T-BoN BO as a valid, eval-efficient, parallel gradient-based Bayesian optimization procedure in the language space.
In addition to its theoretical efficiency, T-BoN BO is simple and easy to implement since it does not require any explicit uncertainty modeling, embeddings, or Gaussian process surrogates. We now demonstrate the empirical performance of our approach using a series of numerical experiments. Concretely, we instantiate T-BoN BO in a digital advertising setting, where the goal is to iteratively refine image-generation prompts for ad creatives so as to maximize ad responsiveness for a population of customers. We test and measure how well and how fast a method optimizes for (i.e., aligns with) the preference distribution induced by LLM-based simulation of personas in the Digital Twins persona dataset called Twin-2k-500 (toubia2025twin). [^2] While LLM-based persona simulation can be biased and may not reflect real-world human behavior (li2025llm; peng2025mega), it still induces a valid preference distribution. As such, it can serve as a valid benchmark on which we can compare the performance of different AI-based optimization methods. We demonstrate that T-BoN BO significantly outperforms strong state-of-the-art baselines such as Best-of-N (snell2024scaling) and GEPA (agrawal2025gepa) in terms of evaluation efficiency, i.e., reaching fixed performance with a smaller number of steps. Specifically, we show that:
  • $\bullet$
    T-BoN BO, the proposed evaluation efficient framework, achieves nearly twice more improvements from Best-of-N baselines compared to GEPA (agrawal2025gepa), the state-of-the-art prompt optimization-based self-improving AI algorithm, which outperforms GRPO (liu2024deepseek), the state-of-the-art reinforcement learning fine-tuning algorithm.
  • $\bullet$
    T-BoN BO outperforms the state-of-the-art baseline even when we simulate preferences without any persona information, i.e., when the evaluator is a single LLM judge with no auxiliary context.
Together, these results show that T-BoN BO can reach good performance both in settings where there is significant heterogeneity in the preferences of the target population as well as in settings with homogeneous preferences.

### v3

# TextBO: Bayesian Optimization in Language Space for Eval-Efficient Self-Improving AI

@@ Abstract
Large Language Models (LLMs) have enabled self-improving AI systems that iteratively generate, evaluate, and refine their outcomes. Recent studies show that prompt-optimization-based self-improvement can outperform state-of-the-art reinforcement-learning fine-tuning of LLMs, but performance is typically measured by generation efficiency. However, in many applications, the constraint is evaluation efficiency: obtaining reliable feedback is far more costly than generating candidates. To optimize for evaluation efficiency, we extend Upper Confidence Bound–Bayesian Optimization (UCB-BO), a framework known for optimal evaluation-efficiency guarantees, to the language domain. Doing so is challenging for two reasons: (i) gradients needed for UCB-BO are ill-defined in discrete prompt space; and (ii) UCB-style exploration relies on a surrogate model and acquisition function, which only live implicitly in the LLM. We overcome these challenges by proving that combining simple textual gradients (LLM-proposed local edits) with the Best-of-N selection strategy statistically emulates ascent along the gradient of the canonical UCB acquisition function. Based on this result, we propose TextBO, a simple, evaluation-efficient self-improving algorithm that operates purely in language space without explicit surrogates or calibrated uncertainty models. We empirically validate TextBO on automated ad-alignment tasks using a persona-induced preference distribution, demonstrating superior performance per evaluation compared to strong baselines such as Best-of-N and GEPA. We also evaluate TextBO’s Best‑of‑N multi‑step textual‑gradient mechanism on agentic AI benchmarks by augmenting GEPA with it and show that it performs better than standard GEPA. In sum, TextBO is a simple and principled framework for AI self-improving system design that bridges prompt optimization with classical Bayesian optimization.
Keywords: Self-improving AI, Evaluation efficiency, Bayesian optimization, Context Engineering, Ad optimization, LLMs, Marketing.
@@ 1 Introduction
1.1 Self-improving AI for Data-driven Decision-making
Data-centric decision making has long relied on an iterative, human-driven cycle of proposing solutions, testing them, and learning from the results. This paradigm underlies many business applications, ranging from website design (hauser2009website) to product recommendations (nandy2021b). We illustrate the cycle using ad optimization (left panel of Figure 1). An ad agency or marketer first generates a slate of creatives from a brief, drawing on product knowledge, theories of consumer behavior, market conditions, and feedback from prior campaigns. This generation step often takes weeks. The candidates are then evaluated (e.g., on advertising platforms or via consumer surveys) to identify the best performers, typically over days to weeks. Finally, the manager analyzes campaign outcomes and uses the insights to guide the next round of creative generation.
This workflow is effective, but it is also constrained by two longstanding limitations. First, iteration speed depends heavily on how quickly humans can generate new candidate solutions, which has historically been the main bottleneck. Second, while evaluation has been studied extensively—with tools such as A/B testing and Multi-Armed Bandits (MAB) (kohavi2020trustworthy; fiez_etal_2024)—there is no comparable formal framework for how practitioners should systematically analyze results and translate them into improved candidates; in practice, this step is often driven by managerial intuition.
Ad generation (Step 1)(Weeks)Analysis (Step 3)(Days)Ad evaluation (Step 2)(Days-weeks) Human-driven cycle BottleneckAd generation (Step 1)(seconds)Analysis (Step 3)(minutes–hours)Ad evaluation (Step 2)(Days-weeks) AI-driven cycle Bottleneck
@@ 1.1 Self-improving AI for Data-driven Decision-making
Recent advances in generative AI have begun to transform this human-driven paradigm by addressing the issues discussed above. Candidate generation is now fast and scalable: in digital advertising, AI models can produce high-quality ad copy, images, and campaign concepts that adhere to brand guidelines in minutes (jansen2024automated; hartmann2025power). [^1] A large share of ad creatives is now AI-generated, reflecting improvements in both quality and speed (coffee2025aiadbuying; Majic2025VibeMarketing). In parallel, LLM-based systems can autonomously conduct analyses by summarizing performance patterns and hypothesizing drivers of success (ferber2024context; guo2024ds; ghosh2024exploring; wiedemer2025videomodelszeroshotlearners). Together, these capabilities make it feasible to automate the full generation–evaluation–analysis loop at scale, allowing systems to update themselves continuously in response to new data (right panel of Figure 1). We refer to such automated, continuously updating systems as self-improving or self-evolving AI (MantiaChatterjeeLee2025; chen2025xbench; silver2025welcome). While there can be many types of self-improving AI systems, a central line of progress has been on iterative prompt optimization, where the AI agent repeatedly refines its own prompts based on performance feedback. Indeed, recent papers show that such approaches can often outperform reinforcement-learning based fine-tuning (agrawal2025gepa); see $\S$2 for a detailed discussion. The key performance objective that this stream of work focuses on is generation efficiency—the number of LLM calls (or candidate solution generations) needed to reach a target performance level.
However, in most business settings, the key bottleneck is a limited evaluation budget rather than a generation budget. For instance, in digital advertising, generating ad creatives is now cheap and fast with modern LLMs/AI models, but ad evaluation, i.e., measuring whether the ad is effective or not, still requires significant human feedback in the form of field deployment (e.g., A/B tests or bandits in ad platforms) or consumer surveys. This is also true for other marketing decisions such as package design, product descriptions, and promotions – in all these cases, generating new candidate solutions is now easy, but evaluating the performance of each solution can take significant time, effort, and human feedback. Beyond business applications, this pattern is also common in many social and scientific domains: AI models like AlphaFold can propose protein structures quickly (jumper2021highly), but their functional properties must still be tested through slow and noisy wet-lab experiment; similarly, AlphaChip can generate chip layouts in silico (goldie2024chip), but performance ultimately depends on fabrication and benchmarking. In sum, the primary goal for self-improving AI systems in business and social settings should be evaluation efficiency. However, as discussed earlier, the existing work has largely focused on generation efficiency as the main objective.
@@ 1.2 Research Agenda and Challenges
In this paper, we seek to design an evaluation-efficient, theoretically grounded, and empirically effective framework for prompt-optimization-based self-improving AI. We define evaluation efficiency as: reaching high-performing solutions using as few costly evaluations as possible.
Our key and novel insight is that iterative prompt-optimization-based self-improving AI can be cast as Bayesian Optimization (BO) in the language space. This recasting then allows us to inherit the key properties of classical BO, in particular, it’s evaluation efficiency. While intuitive and appealing, we nevertheless need to overcome technical challenges in order to translate BO to the language space.
We now briefly summarize canonical BO and discuss the parallels between our problem and canonical BO and provide some intuition for why iterative prompt optimization can be viewed as BO in language space, and the describe the challenges that we need to address to fully leverage the BO framework for our problem.
Bayesian Optimization (BO).
Classical BO is a principled framework for tackling black-box optimization problems with two defining features: (1) There is no closed-form or analytical expression for the objective function or its derivative. In such cases, optimization has to rely on evaluating the function by sampling different points and getting a noisy response. (2) Function evaluation is costly. If evaluation is cheap, then this problem can be solved easily using methods such as grid search or numerical gradient estimation. However, if function evaluation is expensive, then it is important to minimize the number of samples drawn from the black box function. This is the regime in which BO is particularly effective (frazier2018tutorial). In BO, we maintain two models: (1) a surrogate model, that is used for approximating the objective function (given a history of prior evaluations). This model produces a posterior mean and an uncertainty estimate at each point in the space. (2) Acquisition model, that helps guide which point should be chosen for evaluation in the next iteration; this model trades-off predicted value and uncertainty (exploration-exploitation). Among many variants of BO, Upper Confidence Bound (UCB)-BO (srinivas2012information) is known to be evaluation-efficient, i.e., no algorithm can achieve better in terms of both simple regret and expected cumulative regret (whitehouse2023sublinear; wang2023regret). In practice, the UCB acquisition function is often non-concave, so practitioners usually use parallel gradient-based UCB-BO (wilson2018maximizing): start from multiple initial points, take gradient steps to reach several local optima, and evaluate those optima. Parallel gradient-based BO can be especially effective in high-dimensional settings (papenmeier2025understanding).
Iterative prompt optimization as BO.
The iterative-prompt optimization problem can be summarized as follows: in each iteration, the agent repeatedly selects a prompt, observes a noisy performance score, and decides what to try next. The overarching goal is to reach a high-performing prompt (or candidate solution) in as few evaluations as possible. We note that this problem shares some key features of the black-box optimization where BO typically performs well: the prompt-space is high-dimensional and there is no closed-form prompt-to-score model, exploration must be balanced with exploitation when we choose the next prompt to evaluate, and the goal is to maximize progress per costly evaluation. Building on this insight, we propose to cast iterative prompt optimization as parallel gradient-based UCB Bayesian optimization in language space: prompts from prior rounds serve as initializations, gradient-style updates move toward local optima of an implicit UCB acquisition, and in-context learning plays the role of posterior updating.
Challenges.
@@
However, realizing this idea requires overcoming two key challenges. First, in prompt space, gradients are ill-defined because text lacks the inner product structure of a Hilbert space required to measure the directions and magnitudes of change (young1988introduction). Second, even if we can define a meaningful “textual gradient,” we do not have direct access to the UCB acquisition function (or its derivatives). The surrogate posterior over prompt quality that induces the UCB lives implicitly in the LLM, rather than as an explicit Gaussian-process-style model that exposes closed-form posterior means, variances, and their derivatives. Consequently, standard gradient-based BO techniques that differentiate the UCB constructed from a known surrogate cannot be applied naively: we must design a mechanism that elicits UCB-like, optimism-in-the-face-of-uncertainty behavior using only black-box text interactions and critic feedback, without ever computing the UCB function or its gradient in closed form.
1.3 TextBO: Algorithm, Theory, and Numerical Experiments
@@
We propose TextBO, a simple evaluation-efficient prompt-optimization based self-improving AI framework that adapts parallel gradient-based UCB-BO to language space by addressing the above challenges.
To address the first challenge, TextBO builds on TextGrad (textgrad), which adapts gradient descent to discrete text by using an auxiliary LLM critic to propose targeted edits; applying these edits induces a local, directional update (a “textual gradient”) in language space. For example, starting from “A simple photo of a plant-based burger on a white plate,” A critic LLM might suggest: “Make the burger appear delicious, and show it being enjoyed at a barbecue.” Incorporating this edit yields: “A photorealistic image of a juicy plant-based burger with grill marks, served at a vibrant summer barbecue with friends.” Such edits serve as language-native analogues of gradient steps. To address the second challenge, TextBO selects gradient updates via the Best-of-N principle (snell2024scaling), producing what we call the Best-of-N gradient. A single textual gradient step exploits expected improvements implied by past results but does not account for uncertainty. Best-of-N gradient injects exploration into the acquisition by sampling multiple local edits and moving in the direction judged most promising by the critic—an optimism-based rule designed to emulate ascent on an implicit UCB acquisition gradient in language space.
Algorithm.
@@
We now briefly describe our TextBO algorithm. Each iteration in the algorithm consists of three steps. In Step 1 (meta-reflection phase), we use the critic model to analyze the history of previously evaluated solutions and their scores, and update a meta-reflection that summarizes what has tended to work well or poorly. In Step 2 (Best-of-N textual-gradient phase), we perform a sequence of Best-of-N gradient updates from the current prompt. That is, for each gradient step, (i) we propose $N$ textual-gradient edits that transform the current prompt into $N$ candidate prompts, (ii) each candidate prompt is passed to a solution-generation model to produce a corresponding candidate solution, (iii) conditioned on the meta-reflections, the critic model predicts which candidate solution is most likely to perform best, and (iv) we set the prompt associated with this predicted-best solution as the new current prompt. After Step 2, we obtain a refined prompt and its associated solution to evaluate. In Step 3 (evaluation phase), we evaluate this solution to get a scalar score, which TextBO then uses to update its history and meta-reflection for subsequent optimization steps.
Theory.
@@
We provide theoretical guarantees on the performance of our approach. Under mild regularity conditions, we prove that the Best-of-N gradient computed from locally sampled textual edits induces (in probability) an ascent direction of the UCB acquisition function with exploration parameter $\beta_{N}=\Theta(\sqrt{\ln N})$ (Theorem 1). Consequently, TextBO effectively implements parallel, gradient-based UCB Bayesian Optimization in the implicit embedding space induced by the LLM, despite operating purely in language space and without constructing an explicit surrogate or uncertainty model. It therefore inherits the sublinear-regret and evaluation-efficiency guarantees established for UCB-BO, up to constant factors and approximation error due to textual gradients.
Numerical experiments.
We consider two sets of numerical experiments to empirically demonstrate the performance of TextBO – (1) a advertising setting where the goal is to optimize ad creatives, and (2) a series of agentic AI benchmarks from the recent self-improving AI literature.
In the ad optimization experiments, the goal is to iteratively refine image-generation prompts for ad creatives to maximize responsiveness for a population of customers. We test and measure how well and how fast a method optimizes for (i.e., aligns with) the preference distribution induced by LLM-based simulation of personas in the Digital Twins persona dataset called Twin-2k-500 (toubia2025twin). We consider eight advertising scenarios and compare the performance of TextBO to two strong baselines – (1) Best-of-N (snell2024scaling), a popular and well-performing test-time alignment method, and (2) GEPA (agrawal2025gepa), a state-of-the-art prompt optimization based self-improving AI algorithm, which has been shown to outperform reinforcement learning fine-tuning algorithms such as GRPO (liu2024deepseek).
We show that TextBO significantly outperforms both these baselines, across all eight scenarios. Further, in an ablation study, we show that these gains are not simply due to the persona data – even when we simulate preferences without any persona information, i.e., when the evaluator is a single LLM judge with no auxiliary context – TextBO outperforms baseline approaches. Our numerical experiments show that TextBO can reach good performance both in settings where there is significant heterogeneity in the preferences of the target population as well as in settings with homogeneous preferences.
Next, we consider a series of agentic AI benchmark experiments in order to assess whether the same Best‑of‑N multi‑step update principle improves evaluation efficienc. Specifically, we consider the experiments used to evaluate GEPA in agrawal2025gepa – HotpotQA, HoVer, and Pupa. We then compare the performance of TextBO‑GEPA (GEPA augmented with Best‑of‑N gradient sampling and multiple gradient steps) to GEPA. Across all experiments, we show that TextBO-GEPA consistently outperforms GEPA.
Together, these experiments establish TextBO as an effective and general purpose self-improving AI algorithm that can be used for data driven decision-making for both business decisions such as ad creatives, product descriptions, as well as agentic tasks.
1.4 Contributions